OpenStudy (anonymous):

find an equation for the line with points (-5,5) & (5,-4)

OpenStudy (addonexus):

This is undefined. -9/0 is not possible. if it were 0/-9 it would just be 0.

OpenStudy (anonymous):

first get the slope.. m= y2 - y1 / x2 - x1

OpenStudy (anonymous):

i dont have an option for underfined

OpenStudy (anonymous):

so -4 - 5/ 5+ 5

OpenStudy (anonymous):

so your m is -9/10..

OpenStudy (addonexus):

oh wow my mistake

OpenStudy (anonymous):

so y=-9/10 and how do i find the b

OpenStudy (anonymous):

then the slope intercept form. y- y1 = m (x- x1) y-5 = -9/10 (x+ 5)

OpenStudy (anonymous):

i use another method for the equation of the line. b's kinda annoying for me hahaha

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

no its your slope. its not the y

OpenStudy (anonymous):

i know that but isnt that the beginning of the equation?

OpenStudy (anonymous):

multiply both sides by 10.. so its 10y-50 = -9x-45

OpenStudy (anonymous):

yes it is. but depends on you on how you deal with this equation of the line. i usually use this method.

OpenStudy (anonymous):

y=-9/10x+1/2

OpenStudy (anonymous):

10y+ 9x - 5 = 0 thats it

OpenStudy (anonymous):

yes! that can be

OpenStudy (anonymous):

that could also be